OpenStudy (anonymous):

can someone help me with this integral ? ∫6x^3/(2x^4+1)

OpenStudy (amistre64):

the top is almost a derivative of the bottom ... what is the derivative of the bottom?

OpenStudy (amistre64):

also, recall derivative of ln(u) to see why thats one of the first things you try to look at.

OpenStudy (anonymous):

i don't get where the 3/4 comes from...

OpenStudy (amistre64):

what is the derivative of the bottom?

OpenStudy (amistre64):

what is the deriative of: 2x^4 + 1

OpenStudy (anonymous):

8x^3

OpenStudy (amistre64):

good, then we would be sitting pretty if the top was 8x^3, but it aint, so we need to use some math to transform it into an 8

OpenStudy (amistre64):

do you agree that 8/8 = 1?

OpenStudy (anonymous):

lol yes

OpenStudy (amistre64):

\[\int\frac{6x^3}{2x^4+1}\] pull out the 6 and multiply it all by 8/8, why? \[6\int\frac{8}{8}\frac{x^3}{2x^4+1}\] use the top 8 and pull the bottom 8 out of the way \[\frac 68\int\frac{8x^3}{2x^4+1}\] etc ...

OpenStudy (amistre64):

now we integrate it up into: 6/8 ln(2x^4+1) + C

OpenStudy (anonymous):

awesome! thank you!

OpenStudy (amistre64):

youre welcome